Starting from a distance of 20 meters to the left of the origin and at a velocity of 10 m/s, an object accelerates to the right of the origin for 5 seconds at 4 m/s2. What is the position of the object at the end of the 5 seconds of acceleration?

Grade:6

1 Answers

Kushal Chaudhari
5 years ago
In this problem, we may consider that the direction of the object is the positive direction and the initial position x0 = -20 meters (to the left of the origin), the initial velocity u = 10 m/s, the acceleration a = 4 m/s2 and the time is t = 5 seconds. The position is given by 
 
x = (1/2) a t2 + u t + x0
 
= 0.5 * 4 * (5)2 + 10 * 5 - 20 = 80 meters to the right of the origin